the relationship between temperatures in Celsius, C, and in Fahrenheit, F is given F=9/5C+32. If an area has a record low temperature of -30C what is the corresponding Fahrenheit?
Found 3 solutions by lynnlo, Alan3354, ikleyn:
Answer by lynnlo(4176)
Answer by Alan3354(69443)

the relationship between temperatures in Celsius, C, and in Fahrenheit, F is given F=9/5C+32. If an area has a record low temperature of -30C what is the corresponding Fahrenheit?
--------------
-30C = ((-30)*9 + 160)/5F
= (-270 + 160)/5 F
= -110/5 F
= -22 F

Answer by ikleyn(53765)

.
the relationship between temperatures in Celsius, C, and in Fahrenheit, F is given F=9/5C+32.
If an area has a record low temperature of -30C what is the corresponding Fahrenheit?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


Substitute -30 into the formula instead of 'C' and calculate.

Or you prefer others do your job ?